Title

Author’s name (or authors’ names)

Abstract

Put the abstract here.

1 Section heading

1.1 Subsection heading

1.1.1 Subsubsection heading

Example 1.1 Here is an example.

Definition 1.2 Here is a definition.

Lemma 1.3 Here is a lemma.

Theorem 1.4 Here is a theorem.

Proof Proof using the proof environment. □

Conjecture 1.5 Here is a conjecture.

Remark Here is a remark.

The full list of environments defined by the style file is: blank - a new numbered paragraph, claim - Claim, cond - Conditions, conj - Conjecture cor - Corollary, Def - Definition, diag - Diagram, lemma - Lemma prop - Proposition, qn - Question, result - Result, and theorem - Theorem.

Also there are unnumbered environments: aside - Comment, notn - Notation, proof - Proof and rem - Remark.

Some maths commands are already defined, including R, Z and N for the reals, integers and natural numbers.

Please use italics when defining a word for the first time.

Acknowledgements

Put any acknowledgements here.

References

[1] names of author(s), paper title, journal title volume (year), 2000–3000.

[2] name of author(s) (should be A. N. Onymous format), article title, journal title book title, in proceedings title (ed. editor name, single editor), series name for book series, publisher, location of publisher various extra details, volume number for journals (year of journal), (for year of book). (for year of proceedings), page numbers. pp. page numbers in a proceedings.

First address line Second address line Third address line ...

a.b@x.y.z